Last year a real estate agent earned an average of 40,250.00 month. suppose you recently selected a random sample of 25 real estate agents you determined how much each of them earns each month your Computations of their earning resulted to an average of 40,400.00 w/ a standard deviation of 225.00 using a 0.01 Level of significance can it be concluded that the average monthly earnings of estate agents has increased? Assume normality in the population.

1
Expert's answer
2022-01-03T10:19:24-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu\leq 40250"

"H_1:\mu>40250"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=25-1=24" degrees of freedom, and the critical value for a right-tailed test is "t_c = 2.492159."

The rejection region for this right-tailed test is "R = \{t: t > 2.492159\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{40400-40250}{225/\sqrt{25}}\approx3.333333"

Since it is observed that "t = 3.333333 >2.492159= t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed "df=24" degrees of freedom, "t=3.333333" is "p= 0.001388," and since "p = 0.001388 < 0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than "40250," at the "\alpha = 0.01" significance level.

Therefore, there is enough evidence to claim that the average monthly earnings of estate agents has increased, at the "\alpha = 0.01" significance level.